### v1

@@ 4 Heterogeneity and Aggregation Along the Boundary
Our review of the literature indicates that most empirical researchers using BD designs employ pooling-based methods, thereby focusing on (density-weighted) average treatment effects along the entire (or a few segments of the) boundary $\mathcal{B}$. In most implementations, the estimation begins by calculating the distance to the nearest boundary point for every observation, keeping only observations for which this distance is less than a bandwidth, and then continues by pooling all observations in a single local regression analysis, mimicking standard univariate RD methods.
Although the focus on a single scalar parameter is a convenient way of summarizing the information provided by the BD design, this approach fails to exploit its full richness. In this section, we discuss a more general approach for the analysis of BD designs that starts by studying the average treatment effect at each point on the boundary, and then uses this functional causal parameter as the building block for aggregation along the boundary, hence recovering other causal treatment effects. This section summarizes the recent results on identification, estimation and inference discussed in Cattaneo-Titiunik-Yu_2025_BDD-Distance; Cattaneo-Titiunik-Yu_2025_BDD-Location.
@@
which was already implicitly introduced as part of the probability limit emerging from the pooling approaches. However, the functional causal parameter $\tau(\mathbf{x})$ is of independent interest because it captures the average treatment effects at each point on the boundary. Despite capturing potentially interesting causal evidence of heterogeneous treatment effects along the assignment boundary, only a handful of empirical papers have investigated the BATEC or variations thereof. Exceptions include Keele-etal_2017_AIE, velez2019tuning, and gonzalez2021cell.
Another feature of the BATEC is that it provides the key building block for constructing other causal parameters of interest by aggregation along the boundary. For example, Cattaneo-Titiunik-Yu_2025_BDD-Location discuss the following two parameters.
@@
    where $w(\mathbf{x})$ denotes a user-chosen weighing scheme.
@@
The WBATE parameter represents a weighted average of the (potentially heterogeneous) treatment effects $\tau(\mathbf{x})$ at each boundary point. In particular, setting $w(\mathbf{x})=f(\mathbf{x})$ recovers the BATE parameter $\tau$. In practice, other weighting schemes may also be of interest. See Reardon-Robinson_2012_JREE and Wong-Steiner-Cook_2013_JEBS for early methodological discussions, and Rischard-Branson-Miratrix-Bornn_2021_JASA for other examples considered in the context of Bayesian methods for BD designs. The insightful discussion of Qiu-Stoye_2026_BookCh--Discussion provides another application of the WBATE parameter. The LBATE parameter captures the “best” causal treatment effect along the assignment boundary, and thus can be useful for evaluating or developing targeted policies. See, for example, Andrews-Kitagawa-McCloskey_2024_QJE.
Figure 7 gives a graphical representation of the BATEC, WBATE, and LBATE in the context of the same stylized BD design used in Section 3. Because $\tau(\mathbf{x})$ is a function, estimation and inference proceeds by discretizing the boundary to construct local regression estimates for each point $\mathbf{x}\in\mathcal{B}$ on the grid of evaluation points chosen. Figure 7(a) demonstrates the idea with $40$ evenly-spaced grid points, denoted by $\mathbf{b}_{1},\ldots,\mathbf{b}_{40}$. As discussed in the upcoming subsections, distance-based and location-based methods employ these points to conduct estimation and inference, both pointwise and uniformly along the assignment boundary.
Figure 7(b) illustrates graphically the three target causal parameters $(\tau(\mathbf{x}):\mathbf{x}\in\mathcal{B})$, $\tau_{\mathtt{WBATE}}$, and $\tau_{\mathtt{LBATE}}$. The functional causal parameter $\tau(\mathbf{x})$ indicates the presence of heterogeneity along a region of the boundary, while $\tau_{\mathtt{WBATE}}$ and $\tau_{\mathtt{LBATE}}$ give two distinct notions of aggregation of those heterogeneous treatment effects. This stylized example is motivated by the SPP study by LondonoVelezRodriguezSanchez_2020_AEJ introduced above, where $\mathbf{X}_{i}=(X_{1i},X_{2i})=(\text{{SABER11}}_{i},\text{{SISBEN}}_{i})$. Re-analyzing the SPP application, Cattaneo-Titiunik-Yu_2025_BDD-Location found little evidence of heterogeneous treatment effects along the $X_{1i}$ dimension, but declining treatment effects along the $X_{2i}$ dimension. This empirical finding indicates that the average treatment effects along the boundary are roughly similar for all students with low academic performance ($\text{{SABER11}}_{i}=0$) and regardless of their wealth level ($\text{{SISBEN}}_{i}\geq 0$), while heterogeneity in treatment effects is present for the wealthier eligible students ($\text{{SISBEN}}_{i}=0$) as their academic performance increases ($\text{{SABER11}}_{i}\geq 0$).
@@
The discussion so far has focused on causal parameters defined along the entire $\mathcal{B}$, but in some applications researchers have considered different segments of $\mathcal{B}$. Specification (8), and its motivation from the implementation used by LondonoVelezRodriguezSanchez_2020_AEJ and others, gives one example. Employing the BATEC parameter as a building block, it is natural to define causal parameters as integrals or suprema of $\tau(\mathbf{x})$ over a specific region of $\mathcal{B}$. All the methods developed in Cattaneo-Titiunik-Yu_2025_BDD-Distance; Cattaneo-Titiunik-Yu_2025_BDD-Location immediately apply to those cases, offering estimation and inference methods for segment-specific causal parameters based on the core BATEC parameter.
@@ 4.1 Distance-Based Methods
for each unit $i=1,\ldots,n$, and then use this distance to construct a tubular covering $\mathcal{B}$ to implement pooling methods. One example of this approach is ehrlich2018persistent. While the resulting tubular will often be geometrically different from $\mathcal{T}(h)=\{\mathbf{x}\in\mathcal{X}:\mathcal{d}(\mathbf{x},\mathcal{B})\leq h\}$, all the ideas and results discussed in Section 3 continue to apply.
However, the intermediate boundary-point-specific signed distance $(D_{i}(\mathbf{x}):\mathbf{x}\in\mathcal{B})$ for each unit can also be used directly to estimate and conduct inference for the BATEC. The idea is to view the outcomes and the boundary-point-specific scalar distance variables, $(Y_{1},D_{1}(\mathbf{x})),\ldots,(Y_{n},D_{n}(\mathbf{x}))$, as a one dimensional RD design for each point $\mathbf{x}\in\mathcal{B}$. For example, this approach was used by Keele-Titiunik_2015_PA and velez2019tuning to study the effect of TV exposure on voter turnout, using reception and media market boundaries that determine whether citizens are exposed to specific television programming. In these examples, each unit’s score $\mathbf{X}_{i}$ is a latitude-longitude pair that determines their place of residence, the outcome $Y_{i}$ is whether the individual turned out to vote, and the boundary $\mathcal{B}$ is the media market or reception boundary that separates the treated from the control region. In Keele-Titiunik_2015_PA, the treatment is exposure to presidential candidate television ads during the presidential election, and in velez2019tuning the treatment is exposure to Spanish-language television. In both cases, researchers first selected a grid $\mathbf{b}_{1},\mathbf{b}_{2},\ldots\mathbf{b}_{J}$ on the boundary (e.g. Figure 7(a)), and then used the local regression specification:
@@
for each $j=1,\ldots,J$. The key difference is that estimation is conducted with localization for each evaluation point $\mathbf{b}_{j}$ on the boundary, as opposed to localization to the entire boundary as in specifications (1) through (8). Therefore, the estimated coefficient accompanying $T_{i}$ gives a point estimate of $\tau(\mathbf{b}_{j})$ for each $j=1,\ldots,J$
Cattaneo-Titiunik-Yu_2025_BDD-Distance study the large sample properties of the distance-based estimation procedure for BATEC obtained from specification (10). They give necessary and/or sufficient conditions for identification, estimation, and inference in large samples, both pointwise and uniformly along the boundary. Their identification result requires minimal regularity of the boundary $\mathcal{B}$, the distance function $\mathcal{d}(\cdot)$, and the underlying data generating process, building on continuity-based identification in standard univariate RD designs (Hahn-Todd-vanderKlaauw_2001_ECMA). Cattaneo-Titiunik-Yu_2025_BDD-Distance also show that the bias of the distance-based point estimator will depend on the smoothness of the assignment boundary, while the variance of the estimator will be of order $(nh^{2})^{-1}$. Putting these two results together, it follows that bandwidth selection methods for standard univariate RD designs will either minimize the mean square error of the estimator or deliver undersmoothing, depending on the geometry of the assignment boundary and other assumptions. For inference, the authors establish asymptotic validity of confidence intervals and confidence bands, provided the bias of the distance-based point estimator is small enough (which, in turn, crucially depends on the geometry of $\mathcal{B}$).
@@ 4.2 Location-Based Methods
While the distance-based estimator constructed using specification (10) provides a natural bridge between the pooling methods and heterogeneity analysis, Cattaneo-Titiunik-Yu_2025_BDD-Distance demonstrate that this approach would require a smooth assignment boundary for higher-order bias reduction. Otherwise, the distance-based approach will generate a consistent estimator with possibly a high bias, and therefore require a small bandwidth for estimation and inference validity. A solution to this problem is to use local bivariate regression methods.
Cattaneo-Titiunik-Yu_2025_BDD-Location study the properties of local regression treatment effect estimators directly employing the location information encoded in $\mathbf{X}_{i}$ for each unit. The authors develop pointwise and uniform estimation and inference methods for both the BATEC and transformations thereof, such as the WBATE and LBATE. Their local regression specification takes the form:
@@
In the location-based setting, identification of $\tau(\mathbf{x})$ for $\mathbf{x}\in\mathcal{B}$ follows directly from a generalization of continuity-based identification results in the standard univariate RD design (see Hahn-Todd-vanderKlaauw_2001_ECMA; Papay-Willett-Murnane_2011_JoE; Reardon-Robinson_2012_JREE; Wong-Steiner-Cook_2013_JEBS; Keele-Titiunik_2015_PA; Cattaneo-Keele-Titiunik-VazquezBare_2016_JOP). Under minimal regularity conditions on the boundary and data generating process, Cattaneo-Titiunik-Yu_2025_BDD-Location establishes (pointwise and) uniform convergence rates of the form $\sup_{\mathbf{x}\in\mathcal{B}}|\widehat{\tau}(\mathbf{x})-\tau(\mathbf{x})|^{2}=O_{\mathbb{P}}(\tfrac{\log(n)}{nh^{2}}+h^{2(p+1)})$, demonstrating that the location-based estimator does not require stringent smoothness restrictions on the boundary $\mathcal{B}$ to achieve higher order debiasing. Furthermore, the authors establish pointwise and integrated mean square error expansions for the estimator, and develop MSE-optimal bandwidth selection methods. By combining these results, they develop optimal point estimation of the BATEC, both pointwise and uniform, based on specification (11).
For uncertainty quantification, Cattaneo-Titiunik-Yu_2025_BDD-Location consider confidence intervals and confidence bands of the form:
@@
where $\widehat{\mathbb{V}}[\widehat{\tau}(\mathbf{x})]$ denotes any of the usual variance estimators from (weighted) least squares regression methods, and $\mathfrak{c}_{\alpha}$ is the quantile used depending on inferential goal. For confidence intervals, the usual Gaussian quantile is asymptotically valid (i.e., $\mathfrak{c}_{\alpha}=\Phi^{-1}(1-\alpha/2)$), while for confidence bands, a new (larger) quantile needs to be constructed to capture the joint uncertainty features of the estimators $\widehat{\tau}(\mathbf{x})$ for different values along the boundary $\mathcal{B}$. As in the case of pooling-based and distance-based methods, a key condition needed for validity of these confidence intervals and bands is the small (misspecification) bias property. Cattaneo-Titiunik-Yu_2025_BDD-Location address this issue by relying on robust bias correction (Calonico-Cattaneo-Titiunik_2014_ECMA; Calonico-Cattaneo-Farrell_2022_Bernoulli), a construction that proceeds in two steps.
  • Step 1. For a choice $p$ (usually $p=1$), the MSE-optimal bandwidth is computed, and then for that choice of bandwidth the point estimator $\widehat{\tau}(\mathbf{x})$ is computed using specification (11). As a result, $\widehat{\tau}(\mathbf{x})$ is an asymptotically MSE-optimal point estimator of $\tau(\mathbf{x})$.
  • Step 2. For a choice $q>p$ (usually $q=2$), the point estimator $\widehat{\tau}(\mathbf{x})$ and variance estimator $\widehat{\mathbb{V}}[\widehat{\tau}(\mathbf{x})]$ are computed using specification (11) with the same bandwidth used in Step 1. As a result, inference procedures are robust bias-corrected, and hence asymptotically valid.
Cattaneo-Titiunik-Yu_2025_BDD-Location give technical and computational details, which we omit to conserve space, and because they do not fundamentally change the core methodological ideas.
Finally, $\widehat{\tau}(\mathbf{x})$ can also be used to estimate and conduct inference for WBATE and LBATE via plug-in methods. Cattaneo-Titiunik-Yu_2025_BDD-Location provide sufficient conditions and study the validity of such methods. In particular, the WBATE corresponds to an integral over a submanifold of the nonparametric estimator $\widehat{\tau}(\mathbf{x})$. See chen2025semiparametric for related theoretical results when employing series approximations instead of local polynomial regression as commonly done in RD settings.
The general-purpose R package rd2d implement both distance-based and location-based methods (https://rdpackages.github.io/). See Cattaneo-Titiunik-Yu_2025_rd2d for more discussion. Figure 8 gives a stylized example of the type of results that can be obtained via distance-based and location-based methods, building on the same setup as in Figure 1. These figures are motivated by empirical work using the SPP application (LondonoVelezRodriguezSanchez_2020_AEJ).

### v2

@@ 4 Heterogeneity and Aggregation Along the Boundary
Our review of the literature indicates that most empirical researchers using BD designs employ pooling-based methods, thereby focusing on (density-weighted) average treatment effects along the entire (or a few segments of the) boundary $\mathcal{B}$. In most implementations, the estimation begins by calculating the distance to the nearest boundary point for every observation, keeping only observations for which this distance is less than a chosen bandwidth, and then continues by pooling all observations in a single local regression analysis, mimicking standard univariate RD methods.
Although the focus on a single scalar parameter is a convenient way of summarizing the information provided by the BD design, this approach fails to exploit its full richness. In this section, we discuss a more general approach for the analysis of BD designs that starts by studying the average treatment effect at each point on the boundary, and then uses this functional causal parameter as the building block for aggregation along the boundary, hence recovering other causal treatment effects. This section summarizes the recent results on identification, estimation, and inference discussed in Cattaneo et al. (2025a, b).
@@
which was already implicitly introduced as part of the probability limit emerging from the pooling approaches. However, the functional causal parameter $\tau(\mathbf{x})$ is of independent interest because it captures the average treatment effects at each point on the boundary. Despite capturing potentially interesting causal evidence of heterogeneous treatment effects, only a handful of empirical papers have investigated the BATEC or variations thereof. Exceptions include Keele et al. (2017), Velez and Newman (2019), and Gonzalez (2021).
Another feature of the BATEC is that it provides a key building block for constructing other causal parameters of interest by aggregation along the boundary. For example, Cattaneo et al. (2025b) discuss the following two parameters.
@@
    where $w(\mathbf{x})$ denotes a user-chosen weighting scheme.
@@
The WBATE parameter represents a weighted average of the (potentially heterogeneous) treatment effects $\tau(\mathbf{x})$ at each boundary point. In particular, setting $w(\mathbf{x})=f(\mathbf{x})$ recovers the BATE parameter $\tau$. In practice, other weighting schemes may also be of interest. See Reardon and Robinson (2012) and Wong et al. (2013) for early methodological discussions, and Rischard et al. (2021) for other examples considered in the context of Bayesian methods for BD designs. The insightful discussion of Qiu and Stoye (2026) provides another application of the WBATE parameter. The LBATE parameter captures the “best” causal treatment effect along the assignment boundary, and thus can be useful for evaluating or developing targeted policies. See, for example, Andrews et al. (2024).
Figure 7 gives a graphical representation of the BATEC, WBATE, and LBATE in the context of the same stylized BD design used in Section 3. Because $\tau(\mathbf{x})$ is a function, estimation and inference proceed by discretizing the boundary to construct local regression estimates for each point $\mathbf{x}\in\mathcal{B}$ on the grid of evaluation points chosen. Figure 7(a) demonstrates the idea with $40$ evenly-spaced grid points, denoted by $\mathbf{b}_{1},\ldots,\mathbf{b}_{40}$. As discussed in the upcoming subsections, distance-based and location-based methods employ these points to conduct estimation and inference, both pointwise and uniformly along the assignment boundary.
Figure 7(b) illustrates graphically the three target causal parameters $(\tau(\mathbf{x}):\mathbf{x}\in\mathcal{B})$, $\tau_{\mathtt{WBATE}}$, and $\tau_{\mathtt{LBATE}}$. The functional causal parameter $\tau(\mathbf{x})$ indicates the presence of heterogeneity along a region of the boundary, while $\tau_{\mathtt{WBATE}}$ and $\tau_{\mathtt{LBATE}}$ give two distinct notions of aggregation of those heterogeneous treatment effects. This stylized example is motivated by the SPP study by Londoño-Vélez et al. (2020) introduced above, where $\mathbf{X}_{i}=(X_{1i},X_{2i})=(\text{{SABER11}}_{i},\text{{SISBEN}}_{i})$. Re-analyzing the SPP application, Cattaneo et al. (2025b) found little evidence of heterogeneous treatment effects along the $X_{1i}$ dimension, but declining treatment effects along the $X_{2i}$ dimension. This empirical finding indicates that the average treatment effects along the boundary are roughly similar for all students with low academic performance ($\text{{SABER11}}_{i}=0$) and regardless of their wealth level ($\text{{SISBEN}}_{i}\geq 0$), while heterogeneity in treatment effects is present for the wealthier eligible students ($\text{{SISBEN}}_{i}=0$) as their academic performance increases ($\text{{SABER11}}_{i}\geq 0$).
@@
The discussion so far has focused on causal parameters defined along the entire $\mathcal{B}$, but in some applications researchers have considered different segments of $\mathcal{B}$. Specification (8), and its motivation from the implementation used by Londoño-Vélez et al. (2020) and others, gives one example. Employing the BATEC parameter as a building block, it is natural to define causal parameters as integrals or suprema of $\tau(\mathbf{x})$ over a specific region of $\mathcal{B}$. All the methods developed in Cattaneo et al. (2025a, b) immediately apply to those cases, offering estimation and inference methods for segment-specific causal parameters based on the core BATEC parameter.
@@ 4.1 Distance-Based Methods
for each unit $i=1,\ldots,n$, and then use this distance to construct a tubular neighborhood covering $\mathcal{B}$ to implement pooling methods. One example of this approach is Ehrlich and Seidel (2018). While the resulting tubular neighborhood will often be geometrically different from $\mathcal{T}(h)=\{\mathbf{x}\in\mathcal{X}:\mathcal{d}(\mathbf{x},\mathcal{B})\leq h\}$, all the ideas and results discussed in Section 3 continue to apply.
However, the intermediate boundary-point-specific signed distance $(D_{i}(\mathbf{x}):\mathbf{x}\in\mathcal{B})$ for each unit can also be used directly to estimate and conduct inference for the BATEC. The idea is to view the outcomes and the boundary-point-specific scalar distance variables, $(Y_{1},D_{1}(\mathbf{x})),\ldots,(Y_{n},D_{n}(\mathbf{x}))$, as a one-dimensional RD design for each point $\mathbf{x}\in\mathcal{B}$. For example, this approach was used by Keele and Titiunik (2015) and Velez and Newman (2019) to study the effect of TV exposure on voter turnout, using reception and media market boundaries that determine whether citizens are exposed to specific television programming. In these examples, each unit’s score $\mathbf{X}_{i}$ is a latitude-longitude pair that determines their place of residence, the outcome $Y_{i}$ is whether the individual turned out to vote, and the boundary $\mathcal{B}$ is the media market or reception boundary that separates the treated from the control region. In Keele and Titiunik (2015), the treatment is exposure to presidential candidate television ads during the presidential election, and in Velez and Newman (2019) the treatment is exposure to Spanish-language television. In both cases, researchers first selected a grid $\mathbf{b}_{1},\mathbf{b}_{2},\ldots\mathbf{b}_{J}$ on the boundary (e.g. Figure 7(a)), and then used the local regression specification:
@@
for each $j=1,\ldots,J$. The key difference is that estimation is conducted with localization for each evaluation point $\mathbf{b}_{j}$ on the boundary, as opposed to localization to the entire boundary as in specifications (1) through (8). Therefore, the estimated coefficient accompanying $T_{i}$ gives a point estimate of $\tau(\mathbf{b}_{j})$ for each $j=1,\ldots,J$.
Cattaneo et al. (2025a) study the large sample properties of the distance-based estimation procedure for BATEC obtained from specification (10). They give necessary and/or sufficient conditions for identification, estimation, and inference in large samples, both pointwise and uniformly along the boundary. Their identification result requires minimal regularity of the boundary $\mathcal{B}$, the distance function $\mathcal{d}(\cdot)$, and the underlying data generating process, building on continuity-based identification in standard univariate RD designs (Hahn et al., 2001). Cattaneo et al. (2025a) also show that the bias of the distance-based point estimator will depend on the smoothness of the assignment boundary, while the variance of the estimator will be of order $(nh^{2})^{-1}$. Putting these two results together, it follows that bandwidth selection methods for standard univariate RD designs will either minimize the mean square error of the estimator or deliver undersmoothing, depending on the geometry of the assignment boundary and other assumptions. For inference, the authors establish asymptotic validity of confidence intervals and confidence bands, provided the bias of the distance-based point estimator is small enough (which, in turn, crucially depends on the geometry of $\mathcal{B}$).
@@ 4.2 Location-Based Methods
While the distance-based estimator constructed using specification (10) provides a natural bridge between the pooling methods and heterogeneity analysis, Cattaneo et al. (2025a) demonstrate that this approach would require a smooth assignment boundary for higher-order bias reduction. Otherwise, the distance-based approach will generate a consistent estimator with possibly a high bias, and therefore require a small bandwidth for estimation and inference validity. A solution to this problem is to use local bivariate regression methods.
Cattaneo et al. (2025b) study the properties of local regression treatment effect estimators directly employing the location information encoded in $\mathbf{X}_{i}$ for each unit. The authors develop pointwise and uniform estimation and inference methods for both the BATEC and transformations thereof, such as the WBATE and LBATE. Their local regression specification takes the form:
@@
In the location-based setting, identification of $\tau(\mathbf{x})$ for $\mathbf{x}\in\mathcal{B}$ follows directly from a generalization of continuity-based identification results in the standard univariate RD design (see Hahn et al., 2001; Papay et al., 2011; Reardon and Robinson, 2012; Wong et al., 2013; Keele and Titiunik, 2015; Cattaneo et al., 2016b). Under minimal regularity conditions on the boundary and data generating process, Cattaneo et al. (2025b) establishes (pointwise and) uniform convergence rates of the form $\sup_{\mathbf{x}\in\mathcal{B}}|\widehat{\tau}(\mathbf{x})-\tau(\mathbf{x})|^{2}=O_{\mathbb{P}}(\tfrac{\log(n)}{nh^{2}}+h^{2(p+1)})$, demonstrating that the location-based estimator does not require stringent smoothness restrictions on the boundary $\mathcal{B}$ to achieve higher order debiasing. Furthermore, the authors establish pointwise and integrated mean square error expansions for the estimator, and develop MSE-optimal bandwidth selection methods. By combining these results, they develop optimal point estimation of the BATEC, both pointwise and uniform, based on specification (11).
For uncertainty quantification, Cattaneo et al. (2025b) derive confidence intervals and confidence bands of the form:
@@
where $\widehat{\mathbb{V}}[\widehat{\tau}(\mathbf{x})]$ denotes any of the usual variance estimators from (weighted) least squares regression methods, and $\mathfrak{c}_{\alpha}$ is the quantile used depending on inferential goal. For confidence intervals, the usual Gaussian quantile is asymptotically valid (i.e., $\mathfrak{c}_{\alpha}=\Phi^{-1}(1-\alpha/2)$), while for confidence bands, a new (larger) quantile needs to be constructed to capture the joint uncertainty features of the estimators $\widehat{\tau}(\mathbf{x})$ for different values along the boundary $\mathcal{B}$. As in the case of pooling-based and distance-based methods, a key condition needed for validity of these confidence intervals and bands is the small (misspecification) bias property. Cattaneo et al. (2025b) address this issue by relying on robust bias correction (Calonico et al., 2014, 2022), a construction that proceeds in two steps.
  • Step 1. For a choice $\mathsf{P}$ (usually $\mathsf{P}=1$), the point estimator $\widehat{\tau}(\mathbf{x})$ is computed using specification (11) with $p=\mathsf{P}$ and its corresponding MSE-optimal bandwidth. As a result, $\widehat{\tau}(\mathbf{x})$ is an asymptotically MSE-optimal point estimator of $\tau(\mathbf{x})$.
  • Step 2. For a choice $\mathsf{Q}>\mathsf{P}$ (usually $\mathsf{Q}=2$), the point estimator $\widehat{\tau}(\mathbf{x})$ and variance estimator $\widehat{\mathbb{V}}[\widehat{\tau}(\mathbf{x})]$ are computed using specification (11) with $p=\mathsf{Q}$ and the same bandwidth used in Step 1. These two quantities are used to construct the confidence interval estimator $\widehat{\operatorname{I}}(\mathbf{x},\alpha)$. As a result, inference procedures are robust bias-corrected, and hence asymptotically valid.
Cattaneo et al. (2025b) give further technical and computational details, which we omit to conserve space.
Finally, $\widehat{\tau}(\mathbf{x})$ can also be used to estimate and conduct inference for WBATE and LBATE via plug-in methods. Cattaneo et al. (2025b) provide sufficient conditions and study the validity of such methods. In particular, the WBATE corresponds to an integral over a submanifold of the nonparametric estimator $\widehat{\tau}(\mathbf{x})$. See Chen and Gao (2025) for related theoretical results when employing series approximations instead of local polynomial regression as commonly done in RD settings.
The general-purpose R package rd2d implements both distance-based and location-based methods (https://rdpackages.github.io/), see Cattaneo et al. (2025c) for details. Figure 8 gives a stylized example of the type of results that can be obtained via distance-based and location-based methods, building on the same setup as in Figure 1. These figures are motivated by empirical work using the SPP application (Londoño-Vélez et al., 2020).